DIMINISHED CAGE EFFECT IN p-H_{2}: IR IDENTIFICATION OF INTERMEDIATES IN ADDITION RE-ACTIONS OF CL ATOM WITH UNSATURATED HYDROCARBONS

\underline{YUAN-PERN LEE}, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan; MOHAMMED BAHOU, BARBARA GOLEC, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan.

We report infrared absorption spectra of several free radicals produced upon reaction of Cl atoms with unsaturated hydrocarbons isolated in solid p-H_{2}. Cl atoms were produced by in situ photodissociation of Cl_{2} isolated in solid p-H_{2} at 365 nm. In experiments with the Cl_{2}/C_{6}H_{6}/p-H_{2} matrices, intense absorption features at 617.0, 719.8, 956.0, and 1430.5 cm^{−1} and weaker ones at 577.1, 833.6, 876.8, 833.6, 983.0, 993.5, 1008.0, 1026.4, 1112.5, 1118.5, 1179.0, 1406.5, 1509.4, 2967.2, 3054.3, 3063.4, 3070.9, and 3083.9 cm^{−1} appeared upon irradiation of the matrix at 365 nm and increased in intensity upon subsequent annealing of the matrix at 4.8 K for a few minutes. By comparison of vibrational wavenumbers and deuterium isotopic shifts predicted with the B3PW91 and MPW1PW91/aug-cc-pVTZ methods, these features are readily assigned to the σ-complex of ClC_{6}H_{6} (chlorocyclohexadienyl radical), rather than the previously proposed πcomplex. In experiments with the Cl_{2}/C_{2}H_{2}/p-H_{2} matrices, the 1-chloroethyl radicals (CHClCH_{3}) and chloroethene (C_{2}H_{3}Cl) are identified as the main products of the Cl + C_{2}H_{2} reaction in solid p-H_{2}. The assignments of IR absorption lines at 738.2, 1027.6, 1283.4, 1377.1, 1426.6, 1442.6, and 2861.2 cm^{−1} to 1-chloroethyl radicals are based on comparison of the observed vibrational wavenumbers and ^{13}C- and D-isotopic shifts with those predicted with the B3LYP and MP2/augcc-pVDZ methods. These results indicate that the primary products of the addition reaction Cl + C_{2}H_{2}, the 2-chlorovinyl radicals, are unstable; they react readily with p-H_{2} to form CHClCH_{3} and C_{2}H_{3}Cl. If time permits, other examples such as Cl + 1, 3-butadien and H + C_{6}H_{6} or C_{6}H_{5}Cl will be discussed. These results serve as excellent examples to demonstrate that the diminished cage effect of solid p-H_{2} makes production of free radicals via bimolecular reactions feasible.